OpenStudy (loser66):

Explain me, please. find all positive integers such that \(\phi (n) = 6\)

imqwerty (imqwerty):

ϕ(n)=totient function?

OpenStudy (loser66):

yes, this is what my prof said in class, but don't get the reason why

OpenStudy (loser66):

If \(p\geq 11\) we cannot have p|n , why? since then \(p_i -1 =10 \cancel |\phi (n)\) why do we need \(p_i -1 |\phi (n)\)?

imqwerty (imqwerty):

\[n=p_1^{a_1} \times p_2^{a_2} \times p_3^{a_3}...\]p_i are all distinct primes and a_i>=1 \[\phi(n)=\phi(p_1^{a_1}) \times \phi(p_2^{a_2}) \times \phi(p_3^{a_3}).....\] for any p>=11 phi(p) will be >6 and multiplying it with other terms will give phi(n)>6 so we can pay p cannot be >=11

OpenStudy (loser66):

\(a_i\) can be 0

imqwerty (imqwerty):

no i am assuming that \[n=p_1^{a_1} \times p_2^{a_2} \times p_3^{a_3}...\]^this is actual prime factorization of the number n so a_i can't be 0

OpenStudy (loser66):

Got you. Thanks for it. I will work on it. If I don't understand somewhere, I will ask you then. :)

imqwerty (imqwerty):

ok :) lol i too hav to wrk out the nxt step